OpenStudy (anonymous):

A triangle with a base of 1/4 meter has an area of 8 square meters. What is the height, in meters, of the triangle?

OpenStudy (anonymous):

give me 1 second please

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Do you know the formula for the area of a triangle?

OpenStudy (anonymous):

i got an answer but i'll wait to show you it..

OpenStudy (anonymous):

8 square meters

OpenStudy (yttrium):

\[A = \frac{ bh }{ 2 }\] \[h = \frac{ 2A }{ b }\] \[h = \frac{ 2(8) }{ \frac{ 1 }{ 4 } }\] \[h = (16)(4) = \] h = 64 meters

OpenStudy (anonymous):

well. yea thats what i got

OpenStudy (anonymous):

thank you yttrium :)

OpenStudy (mathstudent55):

Correct, 64 m is the height.